+ 0.6* Y, and I-$900. Assume that the following equations fully describe an economy: C = $400 I) (1.5 pts.) How much is the Marginal Propensity to Consume (MPC) for this economy? 2) (1.5 pts.) How much is the Marginal Propensity to Save (MPS) for this economy? (4 pts.) W hat is the equilibrium level of income?
Answer #1

Answer:

a) Marginal Propensity to consume is defined as the change in Consumption as our income changes by unit amount.

MPC= \frac{\partial C}{\partial Y}

In this case it is MPC= 0.6

b) Marginal propensity to save is the defined as the change in saving as our income changes by unit amount.

MPS= 1-MPC= (1-0.6)= 0.4

c) Equilibrium level of income can be calculated as the

Aggregate \ supply = Aggregate \ demand

which can be rewritten as

Y= C+I

Y= 400 + 0.6Y + 900 \\ \Rightarrow Y-0.6Y = 1300 \\ 0.4Y = 1300 \\ Y*= 3250